ProbabilityHard
Question
Let 0 < P(A) < 1, 0 < P(B) < 1 & P(A ∪ B) = P(A) + P(B) - P(A). P(B), then:
Options
A.P(B/A) = P(B) - P(A)
B.P(AC ∪ BC) = P(AC) + P(BC)
C.P((A ∪ B)C) = P(AC). P(BC)
D.P(A/B) = P(A)
Solution
A & B are independent
P(A ∪ B)c = 1 - P(A ∪ B) = 1 - P(A) - P(B) + P(A ∩ B)
= P
- P(B) + P(A) P(B) = P
- P
P(B)
= P
P
P(A ∪ B)c = 1 - P(A ∪ B) = 1 - P(A) - P(B) + P(A ∩ B)
= P
= P
Create a free account to view solution
View Solution FreeMore Probability Questions
A student appears for tests I, III. The student is successful if he passes either in tests I and III. The probabilities ...Three numbers are chosen at random without replacement from {1, 2, 3,......, 10}. The probability that the minimum of th...A bag contains 4 red and 6 black balls. A ball is drawn at random from the bag, its colour is observed and this ball alo...Bag A contains 9 white and 8 black balls, while bag B contains 6 white and 4 black balls. One ball is randomly picked up...For solving a problem, odds against to A are 4 : 3 and odds in favour to B are 7: 5. The probability that the problem wi...